# GeSubNet: Gene Interaction Inference for Disease Subtype Network Generation

### v1

@@ Abstract
Retrieving gene functional networks from knowledge databases presents a challenge due to the mismatch between disease networks and subtype-specific variations. Current solutions, including statistical and deep learning methods, often fail to effectively integrate gene interaction knowledge from databases or explicitly learn subtype-specific interactions. To address this mismatch, we propose GenSubNet, which learns a unified representation capable of predicting gene interactions while distinguishing between different disease subtypes. Graphs generated by such representations can be considered subtype-specific networks. GenSubNet is a multi-step representation learning framework with three modules: First, a deep generative model learns distinct disease subtypes from patient gene expression profiles. Second, a graph neural network captures representations of prior gene networks from knowledge databases, ensuring accurate physical gene interactions. Finally, we integrate these two representations using an inference loss that leverages graph generation capabilities, conditioned on the patient separation loss, to refine subtype-specific information in the learned representation. GenSubNet consistently outperforms traditional methods, with average improvements of 30.6%, 21.0%, 20.1%, and 56.6% across four graph evaluation metrics, averaged over four cancer datasets. Particularly, we conduct a biological simulation experiment to assess how the behavior of selected genes from over 11,000 candidates affects subtypes or patient distributions. The results show that the generated network has the potential to identify subtype-specific genes with an 83% likelihood of impacting patient distribution shifts. The GenSubNet resource is available: https://anonymous.4open.science/r/GeSubNet/
@@ 1 Introduction
Biological knowledge base such as STRING (Szklarczyk et al., 2023) and KEGG (Kanehisa & Goto, 2000), and web-lab experimental datasets such as gene expression data are crucial for understanding disease-gene association. While the knowledge bases are comprehensive, they often lack specificity for disease subtypes. This work introduces a deep learning method to integrate general knowledge bases with disease-subtype specific experimental data to create more targeted knowledge graphs.
Decades of research have generated extensive disease-gene association data, compiled into various biological knowledge databases (Goh et al., 2007b; Szklarczyk et al., 2023; Kanehisa & Goto, 2000). These databases integrate known and predicted gene interactions, forming gene functional networks that describe how gene behaviors relate to disease processes. They support disease research by interpreting experimental results (Vella et al., 2017), facilitating biomarker discovery, and enabling personalized treatment (Goossens et al., 2015).
@@
While biological researchers have proposed a data generation approach to construct meaningful subtype-specific networks (Zaman et al., 2013), these approaches often require extensive in-lab analyses such as pair-wise gene examination among hundreds to thousands of genes. This paper introduces a novel data-driven approach to address this mismatch, automating the integration of gene expression data and knowledge databases to directly generate gene functional networks for various disease subtypes.
Related Works. Existing methods for generating subtype gene networks can be categorized into two groups: statistical and deep learning-based methods. Statistical methods focus on speeding up gene filtering by mining experimental data. These methods employ similarity metrics to measure the correlation between genes. High correlations, such as co-expressed genes (Zhang & Horvath, 2005), are marked as functional interactions. For example, ARACNe (Margolin et al., 2006) uses mutual information to measure expression similarity and removes indirect links with low similarity. WGCNA (Langfelder & Horvath, 2008) calculates Pearson correlation to support large-scale comparisons, while wTO (Gysi et al., 2018) transforms the correlations into probabilistic measures. However, gene interaction retrieval still prioritizes genes of interest.
A few deep learning methods leverage both knowledge databases and experimental datasets. They form disease networks as graphs and embed gene expression data, containing different patient information, as node embeddings. They set up link prediction and reconstruction using graph neural networks (GNNs). The newly reconstructed graphs can be viewed as specific networks. Representative methods include GAERF (Wu et al., 2021), which learns node features with a graph auto-encoder and then uses a random forest to predict links. CSGNN (Zhao et al., 2021) predicts gene interactions using both a mix-hop aggregator and a self-supervised GNN. LR-GNN (Kang et al., 2022) proposes a dynamic graph method to gradually reconstruct graph structure, mitigating the constraints of prior general disease network information. Recent works focus on improving the accuracy of gene-gene link prediction (Li et al., 2024; Pang et al., 2024). However, their objective is only to reconstruct general disease-gene associations, including irrelevant interactions in disease networks. This approach does not explicitly learn the distinct gene interactions unique to disease subtypes.
@@
Contributions and Novelty. We present a new solution for leveraging distinct subtype information from experimental data, i.e., gene expression profiles, to directly infer Gene interactions specific to disease Subtype Networks. This leads us to GenSubNet, which learns a unified representation that can accurately predict prior gene interactions while being able to distinguish different subtypes of a disease. Graphs generated by such representations can be considered subtype-specific networks. GenSubNet is a multi-step learning framework with independent data representation learning and integration. The first step uses a deep generative model to learn gene expression representations. These representations capture distinct data distributions and can distinguish subtypes in a latent feature space. The second step employs a GNN to learn graph representations of prior gene networks. This step ensures GenSubNet captures true gene-gene functional interactions collected in knowledge databases. Finally, we integrate the two representations, updating graph representations and inferring subtype-specific gene interactions using a reconstruction loss on the gene expression data. Our experiments confirm that GenSubNet can simultaneously generate different subtype networks within a general cancer. The contributions lie in:
@@
  • Proposing automated data integration methodology. GenSubNet is an effective architecture that combines a VQ-VAE and Neo-GNN, achieving average improvements of 30.6%, 21.0%, 20.1%, and 56.6% across three metrics on four cancer datasets. More advanced models can be easily integrated into GenSubNet.
  • Impacting Broad Biological Relevance. We propose impactful biological evaluations and a new metric. The experiments involving 11,327 gene evaluations demonstrate that genes selected by GenSubNet are highly related to specific subtypes. We are the first to conduct a simulated experiment, termed Knock-out (Bergman & Siegal, 2003), to assess how the behavior of genes affects different subtypes. The proposed metric evaluates the reliability of selected gene interactions. The results show that GenSubNet effectively narrows down key genes.
  • Integrated Datasets for Cancer Subtyping. We collect physical cancer-gene networks across four knowledge databases and construct machine-learning-ready datasets for experiments and evaluation. We release our datasets with this paper to support continued investigation. The GenSubNet resource is available at: https://anonymous.4open.science/r/GeSubNet/
@@ 2.1 Background: Cancer subtype
Cancer is a major public health concern with increasing incidence and leading to mortality. The National Cancer Institute (NCI) reports that the high costs of cancer care have been projected to grow to \$246.6 billion by 2030 (COS, 2023). A key driver of these high costs and morbidity is cancer’s inherent heterogeneity. Each cancer type is made up of multiple subtypes, characterized by distinct biochemical mechanisms, requiring specific therapeutic approaches (Balmain et al., 2003). While these subtypes may differ biochemically, they often share similar morphological traits, such as the physical structure and form of the organism (Yang et al., 2023), complicating precise diagnosis and treatment responses. This complexity highlights the need for deeper research into gene networks specific to cancer subtypes. However, as shown in Figure 1, current databases like STRING provide only broad cancer gene networks without distinguishing between subtypes such as Luminal A, Luminal B, and Basal-like in breast cancer. This limitation in specificity creates a gap in effectively targeting treatments based on unique subtype characteristics. Our paper addresses this problem by focusing on advancing research and tools that differentiate these subtypes at a more granular level.
@@ 2.2 Problem Setting
Remark. The function $F(\cdot)$ is designed by existing methods focusing on reconstructing the general graph $\mathcal{G}$. The learned representations only carry information for accurate reconstruction. In contrast, we investigate how to learn a representation from both data sources, one that captures essential information from gene interactions while can distinguish different subtypes. Our investigation is based on the following observation: the onset of complex diseases is typically attributed to changes (e.g., perturbations or disruptions) within a limited subset of genes (Goh et al., 2007a).
Formally, given $\mathbf{X}$ and a knowledge graph $\mathcal{G}$, we have $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{|\mathbf{Y}|}\}=F(\mathbf{X};\mathcal{G})$. We aim to learn a unified representation $\mathbf{Z}$ with two properties: $\rm(\hskip 1.79993pti\hskip 1.79993pt)$ Encode high-quality $\mathbf{Z}$ from gene expression profiles $\mathbf{X}$, that is, any $\boldsymbol{z}^{(m)}$ and $\boldsymbol{x}^{(m)}$ should correspond to the same patient group $y$; $\rm(\hskip 1.79993ptii\hskip 1.79993pt)$ Enable $\mathbf{Z}$ to predict gene interactions in $\mathcal{E}$. For the sub-graphs, we have two expectations:
@@
  • Hypothesis-2. $\mathcal{G}_{y}$ should maintain physical and biological meaningfulness. This is an important metric evaluated in our two evaluations, particularly in the Gene Knockout Simulation in Experiment-II, as detailed in Section 4.
Figure 2: Overview of GeSubNet. GeSubNet consists of three modules. Step 1: Patient-M sets up an unsupervised cancer subtyping task to learn the patient sample representation from the input gene expression data, which can distinguish subtypes. Step 2: Graph-M sets up a link prediction task to train the GNN encoder and decoder, learning the graph representation from the input gene graph and expression data. Step 3: Infer-M uses a novel objective function that integrates sample and graph representations. The output from Patient-M conditions the GNN training in Graph-M, which jointly updates the graph structure for subtype-specific networks.
@@ 3.1 Framework
GenSubNet consists of three modules: patient sample representation learning module (Patient-M), graph representation learning module (Graph-M), and network inference module (Infer-M).
  • Patient-M: This module sets up a cancer subtyping task, aiming to project patient gene expression profiles into a latent representation $\mathbf{Z_{p}}$, which can distinguish subtypes. This is typically an unsupervised learning task (Withnell et al., 2021; Yang et al., 2021b; a; 2023). GenSubNet employs a Vector Quantized-Variational AutoEncoder (VQ-VAE) (Van Den Oord et al., 2017) for two purposes: (i) to model this discriminative latent space using a flexible categorical distribution (Chen et al., 2023a), and (ii) to use the decoder as a key component of Infer-M.
  • Graph-M: This module forms a link prediction task, leveraging both knowledge databases and gene expression data to learn $\mathbf{Z_{g}}$. The goal is to train a well-performed GNN autorencoder, where the encoder learns holistic gene interactions, and the decoder is used to generate new graphs. Since we focus on interactions, GenSubNet employs Neo-GNN (Yun et al., 2021), which combines structural information with node representations to prevent over-smoothing of node features.
  • Infer-M: This module involves a novel objective function that integrates $\mathbf{Z_{p}}$ and $\mathbf{Z_{g}}$. GenSubNet uses the information from Patient-M, the decoder, and the reconstruction loss to optimize the prior knowledge in the gene network, i.e., the GNN encoder, for generating subtype-specific networks.
@@ 3.2 Subtype gene network inference
Gene Expression Representation Learning - Patient-M Given a gene expression dataset $\mathbf{X}\in\mathbb{R}^{M\times N}$, we first encode the gene expression sequence to a low-dimensional embedding $\mathbf{Z}_{e}\in\mathbf{R}^{M\times D}$ through linear layers with ReLU activation function:$\mathbf{Z}_{e}=\text{ReLU}(\text{Linear}(\mathbf{X}))$, where $D$ is the dimension of $\mathbf{Z}_{e}$. We apply a Batch Normalization operation to prevent overfitting the limited patient gene expression samples. The $\mathbf{Z}_{e}$ is then projected along the $D$-axis into a set of feature vectors $\mathbf{Z}_{c}\in\mathbb{R}^{M\times D\times S}$, where $S$ denotes the vector dimension. Then, we project $\mathbf{Z}_{c}$ into a discrete codebook (Van Den Oord et al., 2017; Chen et al., 2023b). This involves encoding each dimension of gene features into a code, resulting in $\mathbf{Z}_{p}$ The codebook consists of $K$ latent vectors $\mathcal{P}_{1:K}$, which defines a $K$-way categorical distribution. The projection is conducted using the nearest neighbor search. Then, a decoder, consisting of linear layers with ReLU activations, reconstructs the original gene expression profiles, $\tilde{\mathbf{X}}\in\mathbb{R}^{M\times N}$.
@@
Subtype Network Inference - Infer-M. This module integrates information from both Patient-M and Graph-M to optimize the prior cancer network and generate subtype-specific networks. We propose an objective function that uses Graph-M’s graph generation capabilities, conditioned on the patient separation loss in Patient-M. GenSubNet follows three independent training phases.
@@ 4.1 Dataset and Preprocessing
Gene network dataset. We collected gene functional network resources corresponding to these four cancer types from four well-used knowledge databases, including KEGG (Kanehisa & Goto, 2000), STRING (Szklarczyk et al., 2015), InterPro (Paysan-Lafosse et al., 2023), and Monarch (Mungall et al., 2017). KEGG (KE), STRING (ST), InterPro (Int), and Monarch (Mona).
@@
#Subtype #Feature #Patient #Node #Edge KE ST Int Mona
@@
Baselines. We collected baselines from both the statistical methods and GNN-based methods. (1) The statistical methods include WGCNA (Langfelder & Horvath, 2008), which identifies modules of highly correlated genes using Pearson correlation; wTO (Gysi et al., 2018), which normalizes correlation by all other correlations and calculates probabilities for each edge in the network; ARACNe (Margolin et al., 2006), which calculates mutual information between pairs of nodes and removes indirect relationships; and LEAP (Specht & Li, 2017), which utilizes pseudotime ordering to infer directional relationships. (2) The GNN-based methods include GAERF (Wu et al., 2021), which learns node features with a graph auto-encoder and a random forest classifier; LR-GNN (Kang et al., 2022), which generates node embeddings with a GCN encoder and applies the propagation rule to create links; and CSGNN (Zhao et al., 2021), which predicts node interactions using a mix-hop aggregator and a self-supervised GNN. More details are provided in Appendix C.

### v3

@@ Abstract
Retrieving gene functional networks from knowledge databases presents a challenge due to the mismatch between disease networks and subtype-specific variations. Current solutions, including statistical and deep learning methods, often fail to effectively integrate gene interaction knowledge from databases or explicitly learn subtype-specific interactions. To address this mismatch, we propose GeSubNet, which learns a unified representation capable of predicting gene interactions while distinguishing between different disease subtypes. Graphs generated by such representations can be considered subtype-specific networks. GeSubNet is a multi-step representation learning framework with three modules: First, a deep generative model learns distinct disease subtypes from patient gene expression profiles. Second, a graph neural network captures representations of prior gene networks from knowledge databases, ensuring accurate physical gene interactions. Finally, we integrate these two representations using an inference loss that leverages graph generation capabilities, conditioned on the patient separation loss, to refine subtype-specific information in the learned representation. GeSubNet consistently outperforms traditional methods, with average improvements of 30.6%, 21.0%, 20.1%, and 56.6% across four graph evaluation metrics, averaged over four cancer datasets. Particularly, we conduct a biological simulation experiment to assess how the behavior of selected genes from over 11,000 candidates affects subtypes or patient distributions. The results show that the generated network has the potential to identify subtype-specific genes with an 83% likelihood of impacting patient distribution shifts.
@@ 1 Introduction
Biological knowledge bases such as STRING (Szklarczyk et al., 2023) and KEGG (Kanehisa et al., 2024), and wet-lab experimental datasets such as gene expression data are crucial for understanding disease-gene association. While the knowledge bases are comprehensive, they often lack specificity for disease subtypes. This work introduces a deep learning method to integrate general knowledge bases with disease-subtype-specific experimental data to create more targeted knowledge graphs.
Decades of research have generated extensive disease-gene association data, compiled into various biological knowledge databases (Goh et al., 2007b; Szklarczyk et al., 2023; Kanehisa & Goto, 2000). These databases integrate known and predicted gene interactions, forming gene functional networks that describe how gene behaviors relate to disease processes. They support disease research by interpreting experimental results (Vella et al., 2017), facilitating biomarker discovery (Yang et al., 2023), and enabling personalized treatment (Goossens et al., 2015).
@@
While biological researchers have proposed a data generation approach to construct meaningful subtype-specific networks (Zaman et al., 2013), these approaches often require extensive in-lab analyses such as pair-wise gene examination among hundreds to thousands of gene candidates. This paper introduces a novel data-driven approach to address this mismatch, automating the integration of gene expression data and knowledge databases to directly generate gene functional networks for various disease subtypes.
Related Work. Existing methods for generating subtype gene networks can be categorized into two groups: statistical and deep learning-based methods. Statistical methods focus on speeding up gene filtering by mining experimental data. These methods employ similarity metrics to measure the correlation between genes. High correlations, such as co-expressed genes (Zhang & Horvath, 2005), are marked as functional interactions. For example, ARACNe (Margolin et al., 2006) uses mutual information to measure expression similarity and removes indirect links with low similarity. WGCNA (Langfelder & Horvath, 2008) calculates Pearson correlation to support large-scale comparisons, while wTO (Gysi et al., 2018) transforms the correlations into probabilistic measures. However, gene interaction retrieval still prioritizes genes of interest.
A few deep learning methods leverage both knowledge databases and experimental datasets. They form disease networks as graphs and embed gene expression data, containing different patient information, as node embeddings. They set up link prediction and reconstruction using graph neural networks (GNNs). The newly reconstructed graphs can be viewed as specific networks. Representative methods include GAERF (Wu et al., 2021a), which learns node features with a graph auto-encoder and then uses a random forest to predict links. CSGNN (Zhao et al., 2021) predicts gene interactions using both a mix-hop aggregator and a self-supervised GNN. LR-GNN (Kang et al., 2022) proposes a dynamic graph method to gradually reconstruct graph structure, mitigating the constraints of prior general disease network information. Recent works focus on improving the accuracy of gene-gene link prediction (Li et al., 2024; Pang et al., 2024). However, their objective is only to reconstruct general disease-gene associations, including irrelevant interactions. This approach does not explicitly learn the distinct gene interactions unique to disease subtypes.
@@
Contributions and Novelty. We present a new solution for leveraging distinct subtype information from experimental data, i.e., gene expression profiles, to directly infer Gene interactions specific to disease Subtype Networks. This leads us to GeSubNet, which learns a unified representation that can accurately predict prior gene interactions while being able to distinguish different subtypes of a disease. Graphs generated by such representations can be considered subtype-specific networks. GeSubNet is a multi-step learning framework with independent data representation learning and integration. The first step uses a deep generative model to learn gene expression representations. These representations capture distinct data distributions and can distinguish subtypes in a latent feature space. The second step employs a GNN to learn graph representations of prior gene networks. This step ensures GeSubNet captures true gene-gene functional interactions collected in knowledge databases. Finally, we integrate the two representations, updating graph representations and inferring subtype-specific gene interactions using a reconstruction loss on the gene expression data. Our experiments confirm that GeSubNet can simultaneously generate different subtype networks within a general cancer. The contributions lie in:
@@
  • Proposing Automated Data Integration Methodology. GeSubNet is an effective architecture that combines a VQ-VAE and Neo-GNN, achieving average improvements of 30.6%, 21.0%, 20.1%, and 56.6% across three metrics on four cancer datasets. More advanced models can be easily integrated into GeSubNet.
  • Impacting Broad Biological Relevance. We propose impactful biological evaluations and a new metric. The experiments involving 11,327 gene evaluations demonstrate that genes selected by GeSubNet are highly related to specific subtypes. We are the first to conduct a simulated experiment, termed Knock-out (Bergman & Siegal, 2003), to assess how the behavior of genes affects different subtypes. The proposed metric evaluates the reliability of selected gene interactions. The results show that GeSubNet effectively narrows down key genes.
  • Integrated Datasets for Cancer Subtyping. We collect physical cancer-gene networks across four knowledge databases and construct machine-learning-ready datasets for experiments and evaluation. We release our datasets with this paper to support continued investigation. The code and data resources are available at: https://github.com/chenzRG/GeSubNet
@@ 2.1 Background: Cancer Subtype
Cancer is a major public health concern with increasing incidence and leading to mortality. The National Cancer Institute (NCI) reports that the high costs of cancer care have been projected to grow to \$246.6 billion by 2030 (COS, 2023). A key driver of these high costs and morbidity is cancer’s inherent heterogeneity. Each cancer type is made up of multiple subtypes, characterized by distinct biochemical mechanisms, requiring specific therapeutic approaches (Balmain et al., 2003). While these subtypes may differ biochemically, they often share similar morphological traits, such as the physical structure and form of the organism (Yang et al., 2023), complicating precise diagnosis and treatment responses. This complexity highlights the need for deeper research into gene networks specific to cancer subtypes. However, as shown in Figure 1, current knowledge bases like STRING provide only broad cancer gene networks without distinguishing between subtypes. This limitation in specificity creates a gap in effectively targeting treatments based on unique subtype characteristics. Our paper addresses this problem by focusing on advancing research and tools that differentiate these subtypes at a more granular level.
@@ 2.2 Problem Setting
Remark. The function $F(\cdot)$ is designed by existing methods focusing on reconstructing the general graph $\mathcal{G}$. The learned representations only carry information for accurate reconstruction. In contrast, we investigate how to learn a representation from both data sources, one that captures essential information from gene interactions while distinguishing different subtypes. Our investigation is based on the following observation: the onset of complex diseases is typically attributed to changes (e.g., perturbations or disruptions) within a limited subset of genes (Goh et al., 2007a).
Formally, given $\mathbf{X}$ and a knowledge graph $\mathcal{G}$, we have $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{|\mathbf{Y}|}\}=F(\mathbf{X};\mathcal{G})$. We aim to learn a unified representation $\mathbf{Z}$ with two properties: $\rm(\hskip 1.79993pti\hskip 1.79993pt)$ encode high-quality $\mathbf{Z}$ from gene expression profiles $\mathbf{X}$, that is, any $\boldsymbol{z}^{(m)}$ and $\boldsymbol{x}^{(m)}$ should correspond to the same patient group $y$; $\rm(\hskip 1.79993ptii\hskip 1.79993pt)$ enable $\mathbf{Z}$ to predict gene-gene interactions in $\mathcal{E}$. For the sub-graphs, we have two hypotheses:
@@
  • Hypothesis-2. $\mathcal{G}_{y}$ must maintain physical and biological meaningfulness. This is an important metric evaluated in our two evaluations, particularly in the Gene Knockout Simulation in Experiment-II, as detailed in Section 4.
Figure 2: Overview of GeSubNet. GeSubNet consists of three modules. Step 1: Patient-M sets up an unsupervised cancer subtyping task to learn the patient sample representation ($\mathbf{Z_{p}}$) from the input gene expression data ($\mathbf{X}$), which can distinguish subtypes. Step 2: Graph-M sets up a link prediction task to train the GNN encoder and decoder, learning the graph representation ($\mathbf{Z_{g}}$) from the input gene graph ($\mathcal{G}$) and expression data ($\mathbf{X}$). Step 3: Infer-M uses an objective function that integrates representations to generate subtype-specific networks. The reconstruction from Patient-M, conditioned on the GNN training in Graph-M ($q_{\theta}(\mathbf{z_{g}}|\mathcal{G})$), refines the graph structure, which can maintain accurate patient profile reconstruction ($p_{\phi}(\mathbf{x}|\mathbf{\tilde{x}})$).
@@ 3.1 Framework
GeSubNet consists of three modules: patient sample representation learning module (Patient-M), graph representation learning module (Graph-M), and network inference module (Infer-M).
  • Patient-M: This module sets up a cancer subtyping task, aiming to project patient gene expression profiles into a latent representation $\mathbf{Z_{p}}$, which can distinguish subtypes. This is typically an unsupervised learning task (Withnell et al., 2021; Yang et al., 2021b; a; 2023). GeSubNet employs a Vector Quantized-Variational AutoEncoder (VQ-VAE) (Van Den Oord et al., 2017) for two purposes: (i) to model this discriminative latent space using a flexible categorical distribution (Chen et al., 2023a), and (ii) to use the decoder as a key component of Infer-M.
  • Graph-M: This module forms a link prediction task, leveraging both knowledge databases and gene expression data to learn $\mathbf{Z_{g}}$. The goal is to train a well-performed GNN autorencoder, where the encoder learns holistic gene interactions, and the decoder is used to generate new graphs. Since we focus on interactions, GeSubNet employs Neo-GNN (Yun et al., 2021), which combines structural information with node representations to prevent over-smoothing of node features.
  • Infer-M: This module involves a novel objective function that integrates $\mathbf{Z_{p}}$ and $\mathbf{Z_{g}}$. GeSubNet uses the information from Patient-M, the decoder, and the reconstruction loss to optimize the prior knowledge in the gene network, i.e., the GNN encoder, for generating subtype-specific networks.
@@ 3.2 Subtype Gene Network Inference
Gene Expression Representation Learning - Patient-M. Given a gene expression dataset $\mathbf{X}\in\mathbb{R}^{M\times N}$, we first encode the gene expression profile to a low-dimensional embedding $\mathbf{Z}_{e}\in\mathbf{R}^{M\times D}$ through linear layers with ReLU activation function: $\mathbf{Z}_{e}=\text{ReLU}(\text{Linear}(\mathbf{X}))$, where $D$ is the dimension of $\mathbf{Z}_{e}$. We apply a Batch Normalization operation to prevent overfitting the limited patient gene expression samples. The $\mathbf{Z}_{e}$ is then projected along the $D$-axis into a set of feature vectors $\mathbf{Z}_{c}\in\mathbb{R}^{M\times D\times S}$, where $S$ denotes the vector dimension. Then, we project $\mathbf{Z}_{c}$ into a discrete codebook (Van Den Oord et al., 2017; Chen et al., 2023b). This involves encoding each dimension of gene features into a code, resulting in $\mathbf{Z}_{p}$ The codebook consists of $K$ latent vectors $\mathcal{P}_{1:K}$, which defines a $K$-way categorical distribution. The projection is conducted using the nearest neighbor search. Then, a decoder, consisting of linear layers with ReLU activations, reconstructs the original gene expression profiles, $\tilde{\mathbf{X}}\in\mathbb{R}^{M\times N}$.
@@
Subtype Network Inference - Infer-M. This module integrates information from both Patient-M and Graph-M to optimize the prior cancer network and generate subtype-specific networks. We propose an objective function that uses Graph-M’s graph generation capabilities, conditioned on the patient separation loss in Patient-M. GeSubNet follows three independent training phases.
@@ 4.1 Dataset and Preprocessing
Gene network dataset. We collected gene functional networks corresponding to these four cancer types from four well-used knowledge databases, including KEGG (KE) (Kanehisa & Goto, 2000), STRING (ST) (Szklarczyk et al., 2015), InterPro (Int) (Paysan-Lafosse et al., 2023), and Monarch (Mona) (Mungall et al., 2017).
@@
#Subtypes #Features #Patients #Nodes #Edges KE ST Int Mona
@@
Baselines. We collected baselines from both the statistical methods and GNN-based methods. (1) The statistical methods include WGCNA (Langfelder & Horvath, 2008), which identifies modules of highly correlated genes using Pearson correlation; wTO (Gysi et al., 2018), which normalizes correlation by all other correlations and calculates probabilities for each edge in the network; ARACNe (Margolin et al., 2006), which calculates mutual information between pairs of nodes and removes indirect relationships; and LEAP (Specht & Li, 2017), which utilizes pseudotime ordering to infer directional relationships. (2) The GNN-based methods include GAERF (Wu et al., 2021a), which learns node features with a graph auto-encoder and a random forest classifier; LR-GNN (Kang et al., 2022), which generates node embeddings with a GCN encoder and applies the propagation rule to create links; and CSGNN (Zhao et al., 2021), which predicts node interactions using a mix-hop aggregator and a self-supervised GNN. More details are provided in Appendix C.